# A Survey on Knowledge Organization Systems of Research Fields: Resources and Challenges

### v1

@@ 3. Sub-challenge 3:
Semantic MediaWiki is an open-source extension of MediaWiki, the engine that runs underneath Wikipedia (Vrandečić and Krötzsch, 2009). Semantic MediaWiki provides an environment that is stable, powerful, and scalable, and it enables users to browse and collaboratively edit ontologies. It also provides facilities for rating pages and users.
@@
In brief, there is a wide collection of tools that can facilitate the integration of the different KOSs. However, all these tools are general-purpose and built to support a multitude of use cases. As a result, users that do not have experience with these technologies and semantic web technologies might find them quite complex and difficult to learn. We still lack user-friendly tools that are able to effectively support the creation and curation of KOSs.
@@ 5.3 Improving the language coverage
English is the de-facto language of Science nowadays (Sugimoto and Lariviére, 2018). Most international conferences are held in English, and the world’s top scientific journals are published in English. However, some countries, like China, Russia, and Japan, have several dedicated journals and conferences publishing scientific papers in their own language. It is crucial then to have KOSs in other languages as well, so as to facilitate interoperability and the cross-language exploration and exploitation of digital artefacts (Lei Zeng and Mai Chan, 2004). This would also benefit students in non-English speaking countries, allowing them to understand and navigate KOSs without requiring fluency in English.
In Section 4.3.6, we acknowledged that some KOSs already have partial translations in other languages, such as the Dewey Decimal Classification, the Agrovoc Thesaurus and the Unesco Thesaurus. However, the process is far from being complete. The Unesco Thesaurus is available only in five languages (English, Arabic, Russian, French, and Spanish). The Agrovoc Thesaurus has a very wide variety of languages; however, not all the forty-two languages are equally represented, as reported in Fig. 3. Notably, the Food and Agriculture Organization, who coordinates Agrovoc, allows international institutions to contribute to Agrovoc by authoring translations. This is an interesting solution that may be beneficial for several other KOSs.
@@ 5.4 Reconciliating expert disagreements
In the literature, we can find different works studying disagreement among experts in the context of building KOSs. Gu et al. (2007) audited the semantic types assigned to UMLS concepts with the support of four experts. In particular, the experts assessed whether the current semantic types assigned to the concepts were correct. In some cases, the disagreement prevailed even after multiple rounds and was eventually resolved with a subsequent discussion. This experience indicates that the process of mitigating disagreement between experts is quite challenging and certainly time-consuming. Osborne et al. (2019) analysed the agreement of experts in characterising articles according to research areas in Software Architecture and highlighted that, in this domain, most of the disagreement was between domain experts of different seniority (e.g., Full Professor vs. PhD students). Fan and Friedman (2007) also explored the issue of disagreement among experts on research topics. They developed a classifier to annotate research documents with concepts derived from UMLS and then asked experts to categorise UMLS concepts into nine different types. While the inter-annotator agreement was relatively high (0.82), their ablation study revealed that experts often disagreed on vague concepts. For instance, the concept of "promotion" was classified both as a behaviour and as a biological function. The authors emphasise that unresolved disagreements can introduce errors into gold standards, which in turn can propagate into downstream applications.
@@ 5.5 Assessing the quality of KOSs
The second case of inconsistency is when there is a transitive hierarchy [^43] Transitive hierarchies - https://www.w3.org/TR/skos-primer/#sectransitivebroader: “A” is broader than “B”, “B” is broader than “C”, but “A” is also broader than “C”. Although, in both cases “A” is broader than “C”, hence it may not be considered as a real mistake, however it adds confusion to the hierarchical structure which necessitates to be addressed. Besides, by convention, skos:narrower and skos:broader must be used to assert a direct or immediate hierarchical link between concepts and are not declared as transitive properties.
@@
The literature presents various approaches for assessing the quality of KOSs (Morrey et al., 2009). For instance, Raad and Cruz (2015) discusses four categories of ontology evaluation approaches: i) gold standard-based, ii) corpus-based, iii) task-based, and iv) criteria based. Gold standard-based evaluation approaches compare the KOS with a reference system. Corpus-based evaluation approaches assess the extent of coverage of the system in a given domain. Task-based evaluation approaches measure to what extent the system helps to improve a task. Finally, criteria-based evaluation approaches measure the extent of a system in adhering to certain criteria. The wide range of evaluation methods highlights the complexity of assessing the quality of a KOS. Indeed, for gold standard-based evaluations, Raad and Cruz (2015) point out that it is hard to find a suitable gold standard because it needs to be created with the same conditions as the evaluated system. Corpus-based evaluation approaches also have similar issues. On the other hand, task-based and criteria-based evaluation approaches have fewer challenges to tackle, but they fall short in assessing KOSs at the conceptual level (i.e., accuracy and completeness).
@@ 5.7 Automatic generation and updating KOSs
The literature presents various automated solutions for creating KOSs. These can either be employed to generate KOSs from scratch, e.g., (Osborne and Motta, 2015), or to scan the recent literature and extend existing ones, e.g., (Huang et al., 2020; Pisu et al., 2024). Klink-2 (Osborne and Motta, 2015) demonstrated high effectiveness in automatically generating large-scale and granular ontologies of research areas. Other approaches include Shang et al. (2020), who introduced NetTaxo, which is an automated topic taxonomy construction framework that leverages both text data and network structures to build hierarchical topic representations. Whereas, Zhang et al. (2018) presented TaxoGen, an unsupervised method that employs term embeddings and hierarchical clustering to recursively build a topic taxonomy. On the other hand, Huang et al. (2020) introduced a technique for building topic taxonomies guided by seed concepts. Given a text corpus and an initial taxonomy of concepts, their approach develops a more comprehensive taxonomy. This approach is potentially valuable in the context of updating existing KOSs.
@@ 6 Threats to validity
With regard to the inclusion and exclusion criteria, we identified two main potential threats to external validity. The first concerns the exclusion of KOSs that do not offer the English version. This was set due to the predominance of English as the language of scientific communication (Sugimoto and Lariviére, 2018). Additionally, the absence of English versions would have hindered our ability to conduct in-depth analyses, such as the one presented in Table 2. The second threat arises from the exclusion of KOSs that are designed specifically for individual digital libraries and have not been widely adopted by the broader community. This exclusion was necessary to avoid an unmanageable increase in the number of KOS candidates, which would have made the analysis impractical. Additionally, these KOSs are often customised to fit the unique content of their respective libraries, which could result in a skewed portrayal of the broader scientific landscape.
@@ 7 Conclusions
Our findings indicate that the current generation of KOSs requires substantial enhancements in scope, quality, and granularity. Notably, seven disciplines (History, Political Science, Environmental Science, Materials Science, Geography, Sociology, and Business) lack dedicated, field-specific KOSs. Additionally, among the existing multidisciplinary KOSs, only five provide comprehensive coverage of the 19 fields analysed in this study.
The analysed KOSs exhibit considerable diversity in both the number of concepts and structural depth. Some KOSs include as few as 48 concepts (e.g., Fields of research and development) with a relatively shallow structure, while others encompass more than 3 million concepts and feature a very deep structure, extending beyond 30 levels (e.g., Open Biological and Biomedical Ontology and Unified Medical Language System). The majority of the KOSs analysed (24) are still traditional taxonomies. However, a growing number of more recent KOS (18) are now based on ontologies. Regarding hierarchical structures, there is a fairly balanced distribution: 24 KOSs utilise a poly-hierarchical structure, while 20 employ a mono-hierarchical one. Traditionally, Knowledge Organization Systems (KOSs) have been created and maintained manually. However, in recent years, there has been a significant shift toward more automated methods. Consequently, eight KOSs, including OpenAlex Topics and Microsoft’s Fields of Study, have implemented modern automatic or semi-automatic pipelines for their generation and updating processes. Eighteen KOSs are updated annually, demonstrating a significant commitment to maintaining up-to-date resources in some disciplines (e.g., “Medicine”, “Engineering”, “Agriculture” and “Computer Science”). Finally, 26 KOSs are available under open licenses.
@@
Overall, this is a promising area of research that has yet to fully capitalise on recent advances in generative AI, which hold significant potential to drive progress in the field. To achieve this, we believe it is essential for the Open Science, Digital Libraries, and Artificial Intelligence communities to collaborate in developing a unified framework of interlinked resources. We hope that this survey paper serves as a meaningful first step in this direction.

### v2

@@ 3. Sub-challenge 3:
Semantic MediaWiki is an open-source extension of MediaWiki, the engine that runs underneath Wikipedia (Vrandečić and Krötzsch, 2009). Semantic MediaWiki provides a stable, powerful, and scalable environment, enabling users to browse and collaboratively edit ontologies. It also provides facilities for rating pages and users.
VocBench3 is an open-source web application that allows users to create, manage, and share ontologies and thesauri (Stellato et al., 2020). It provides a user-friendly interface for editing and managing RDF data, supports collaborative editing and version control, enabling tracking of changes and the ability to revert to prior versions.
@@
In brief, a wide collection of tools facilitating the integration of the different KOSs is available. However, all these tools are general-purpose and built to support a multitude of use cases. As a result, users unfamiliar with the relevant technologies and the semantic web might find them quite complex and difficult to learn. We still lack user-friendly tools that are able to effectively support the creation and curation of KOSs.
@@ 5.3 Improving the language coverage
English is the de-facto language of Science nowadays (Sugimoto and Larivière, 2018). Most international conferences are held in English, and the world’s top scientific journals are published in English. However, some countries, like China, Russia, and Japan, have several dedicated journals and conferences publishing scientific papers in their own language. It is crucial then to have KOSs in other languages as well so as to facilitate interoperability and the cross-language exploration and exploitation of digital artefacts (Lei Zeng and Mai Chan, 2004). This would also benefit students in non-English speaking countries, allowing them to understand and navigate KOSs without requiring fluency in English.
In Section 4.3.6, we acknowledged that some KOSs already have partial translations in other languages, such as the Dewey Decimal Classification, the Agrovoc Thesaurus and the Unesco Thesaurus. However, the process is far from being complete. The Unesco Thesaurus is available only in five languages (English, Arabic, Russian, French, and Spanish). The Agrovoc Thesaurus has a very wide variety of languages; however, not all the forty-two languages are equally represented, as reported in Figure 5. Notably, the Food and Agriculture Organization, who coordinates Agrovoc, allows international institutions to contribute to Agrovoc by authoring translations. This is an interesting solution that may be beneficial for several other KOSs.
@@ 5.4 Reconciliating expert disagreements
In the literature, we can find different works studying disagreement among experts in the context of building KOSs. Gu et al. (2007) audited the semantic types assigned to UMLS concepts with the support of four experts. In particular, the experts assessed whether the current semantic types assigned to the concepts were correct. In some cases, the disagreement prevailed even after multiple rounds and was eventually resolved with a subsequent discussion. This experience indicates that the process of mitigating disagreement between experts is quite challenging and certainly time-consuming. Osborne et al. (2019) analysed the agreement of experts in characterising articles according to research areas in Software Architecture and highlighted that, in this domain, most of the disagreement was between domain experts of different seniority (e.g., Full Professor vs. PhD students). Fan and Friedman (2007) also explored the issue of disagreement among experts on research topics. They developed a classifier to annotate research documents with concepts derived from UMLS and then asked experts to categorise UMLS concepts into nine different types. While the inter-annotator agreement was relatively high (0.82), their ablation study revealed that experts often disagreed on vague concepts. For instance, the concept of “promotion” was classified both as a behaviour and as a biological function. The authors emphasise that unresolved disagreements can introduce errors into gold standards, which in turn can propagate into downstream applications.
@@ 5.5 Assessing the quality of KOSs
The second case of inconsistency is when there is a transitive hierarchy [^45] Transitive hierarchies — https://www.w3.org/TR/skos-primer/#sectransitivebroader: “A” is broader than “B”, “B” is broader than “C”, but “A” is also broader than “C”. Although, in both cases “A” is broader than “C”, hence it may not be considered as a real mistake, however it adds confusion to the hierarchical structure which necessitates to be addressed. Besides, by convention, skos:narrower and skos:broader must be used to assert a direct or immediate hierarchical link between concepts and are not declared as transitive properties.
@@
The literature presents various approaches for assessing the quality of KOSs (Morrey et al., 2009). For instance, Raad and Cruz (2015) discusses four categories of ontology evaluation approaches: i) gold standard-based, ii) corpus-based, iii) task-based, and iv) criteria-based. Gold standard-based evaluation approaches compare the KOS with a reference system. Corpus-based evaluation approaches assess the extent of coverage of the system in a given domain. Task-based evaluation approaches measure to what extent the system helps to improve a task. Finally, criteria-based evaluation approaches measure the extent of a system in adhering to certain criteria. The wide range of evaluation methods highlights the complexity of assessing the quality of a KOS. Indeed, for gold standard-based evaluations, Raad and Cruz (2015) point out that it is hard to find a suitable gold standard because it needs to be created with the same conditions as the evaluated system. Corpus-based evaluation approaches also have similar issues. On the other hand, task-based and criteria-based evaluation approaches have fewer challenges to tackle, but they fall short in assessing KOSs at the conceptual level (i.e., accuracy and completeness).
@@ 5.7 Automatic generation and updating KOSs
The literature presents various automated solutions for creating KOSs. These can either be employed to generate KOSs from scratch, e.g., (Osborne and Motta, 2015), or to scan the recent literature and extend existing ones, e.g., (Huang et al., 2020; Pisu et al., 2024). Klink-2 (Osborne and Motta, 2015) demonstrated high effectiveness in automatically generating large-scale and granular ontologies of research areas. On a similar note, Shang et al. (2020) introduced NetTaxo, which is an automated topic taxonomy construction framework that leverages both text data and network structures to build hierarchical topic representations. Whereas, Zhang et al. (2018) presented TaxoGen, an unsupervised method that employs term embeddings and hierarchical clustering to recursively build a topic taxonomy. On the other hand, Huang et al. (2020) introduced a technique for building topic taxonomies guided by seed concepts. Given a text corpus and an initial taxonomy of concepts, their approach develops a more comprehensive taxonomy. This approach is potentially valuable in the context of updating existing KOSs.
@@ 6 Threats to validity
With regard to the inclusion and exclusion criteria, we identified two main potential threats to external validity. The first concerns the exclusion of KOSs that do not offer the English version. This was set due to the predominance of English as the language of scientific communication (Sugimoto and Larivière, 2018). Additionally, the absence of English versions would have hindered our ability to conduct in-depth analyses, such as the one presented in Table 2. The second threat arises from the exclusion of KOSs that are designed specifically for individual digital libraries and have not been widely adopted by the broader community. This exclusion was necessary to avoid an unmanageable increase in the number of KOS candidates, which would have made the analysis impractical. Additionally, these KOSs are often customised to fit the unique content of their respective libraries, which could result in a skewed portrayal of the broader scientific landscape.
@@ 7 Conclusions
Our findings indicate that the current generation of KOSs requires substantial enhancements in scope, quality, and granularity. Notably, seven disciplines (“History”, “Political Science”, “Environmental Science”, “Material Science”, “Geography”, “Sociology”, and “Business”) lack dedicated, field-specific KOSs. Additionally, among the existing multidisciplinary KOSs, only five provide comprehensive coverage of the 19 fields analysed in this study.
The analysed KOSs exhibit considerable diversity in both the number of concepts and structural depth. Some KOSs include as few as 48 concepts (e.g., Fields of research and development) with a relatively shallow structure, while others encompass more than 3 million concepts and feature a very deep structure, extending beyond 30 levels (e.g., Open Biological and Biomedical Ontology and Unified Medical Language System). The majority of the KOSs analysed (24) are still traditional taxonomies. However, a growing number of more recent KOSs (18) are now based on ontologies. Regarding hierarchical structures, there is a fairly balanced distribution: 24 KOSs utilise a poly-hierarchical structure, while 20 employ a mono-hierarchical one. Traditionally, KOSs have been created and maintained manually. However, in recent years, there has been a significant shift toward more automated methods. Consequently, eight KOSs, including OpenAlex Topics and Microsoft’s Fields of Study, have implemented modern automatic or semi-automatic pipelines for their generation and updating processes. Eighteen KOSs are updated annually, demonstrating a significant commitment to maintaining up-to-date resources in some disciplines (e.g., “Medicine”, “Engineering”, “Agriculture” and “Computer Science”). Finally, 26 KOSs are available under open licenses.
@@
Overall, this is a promising area of research that has yet to fully capitalise on recent advances in generative AI, which hold significant potential to drive progress in the field. To achieve this, we believe it is essential for the Open Science, Digital Libraries, and Artificial Intelligence communities to collaborate in developing a unified framework of interlinked resources. We hope that this survey paper serves as a meaningful first step toward this direction.